# Register Always Matters: Analysis of LLM Pretraining Data Through the Lens of Language Variation

### v1

@@ Abstract
Pretraining data curation is a cornerstone in Large Language Model (LLM) development, leading to growing research on quality filtering of large web corpora. From statistical quality flags to LLM-based labeling systems, datasets are divided into categories, frequently reducing to a binary: those passing the filters deemed as valuable examples, others discarded as useless or detrimental. However, a more detailed understanding of the contribution of different kinds of texts to model performance is still largely lacking. In this article, we present the first study utilizing registers (also known as genres)—a widely used standard in corpus linguistics to model linguistic variation—to curate pretraining datasets and investigate the effect of register on the performance of LLMs. We perform comparative studies by training models with register classified data and evaluating them using standard benchmarks, and show that the register of pretraining data substantially affects model performance. We uncover surprising relationships between the pretraining material and the resulting models: using the News register results in subpar performance, and on the contrary, including the Opinion class, covering texts such as reviews and opinion blogs, is highly beneficial. While a model trained on the entire unfiltered dataset outperforms those trained on datasets limited to a single register, combining well-performing registers like How-to-Instructions, Informational Description, and Opinion leads to major improvements. Furthermore, analysis of individual benchmark results reveals key differences in the strengths and drawbacks of specific register classes as pretraining data: How-to-Instructions excels at physical reasoning and sentence completion while barely crossing random baselines on world-knowledge benchmarks, while Narrative boosts performance on social interaction tasks but struggles with scientific questions. These findings show that register is an important explainer of model variation and can facilitate more deliberate future data selection practices.
@@ 1 Introduction
Recent work on scaling laws and the limits of human-generated training data available for Large Language Model (LLM) training have caused an uptick in research on pretraining data quality as opposed to quantity (Kaplan et al., 2020; Hoffmann et al., 2022; Muennighoff et al., 2023; Villalobos et al., 2024). While state-of-the-art models are still trained predominantly on large web corpora, there is increasing focus on filtering such data to remove potentially detrimental material. The selection of pretraining data often relies on simple heuristics or quality signals (Longpre et al., 2024; Ostendorff et al., 2024), sometimes together with ML-based labels (Weber et al., 2024) and LLM-created quality flags (Henriksson et al., 2025). However, language use varies along a number of dimensions other than the presumed dichotomy of positive or negative value for LLM training—ranging from persuasive, narrative, to informative texts. The effects of this variation on LLM capabilities remain unknown. In this paper, we approach pretraining data curation from this perspective: how do different registers (or genres) of training data affect LLM performance on commonly used benchmarks.
@@
Register studies is a widely applied paradigm in corpus linguistics that examines language variation across different situations, from casual spoken conversations to persuasive and informational settings (Biber, 1988; 1995; Biber & Conrad, 2019). Registers are defined as situationally characterized text varieties, with typical classes encompassing news, reviews and song lyrics. Register has been shown to be one of the most important predictors of linguistic variation (Biber, 2012)—and, as put by the founder of register studies, Douglas Biber, “register always matters(Biber, 2013).
@@
Web register classification has long aimed to characterize different kinds of web texts in order to better utilize and understand the web as a corpus (Kilgarriff & Grefenstette, 2003; Baroni & Kilgarriff, 2006). However, early studies focused on pre-selected sets of registers presumed to be typical of the web, and thus lacked generalizability (Santini, 2007; Sharoff et al., 2010) and subsequently failed to capture the full range of linguistic variation found on the unrestricted web. Egbert et al. (2015) addressed this gap by creating a corpus representing the full spectrum of English web registers, with Laippala et al. (2023) presenting the first deep learning-based web register classifier trained on the data and targeting the entire unrestricted web. Other studies expanded this approach to multilingual settings (Repo et al., 2021; Rönnqvist et al., 2021; Kuzman et al., 2023a; Henriksson et al., 2024), showing that web registers are identifiable in web-scale data. These advances now enable the integration of corpus linguistics frameworks with LLM pretraining data curation—our study being to the best of our knowledge the first to bridge these areas, despite the well-documented role of registers in linguistic variation.
@@
To understand the characteristics of web registers in the context of LLM pretraining, we train small generative models up to 100 billion tokens with web data filtered to individual registers. We evaluate these models using well-known benchmarks and analyze the results through average accuracy and performance on individual benchmarks, revealing how each register impacts model capabilities. Our results show that register has a substantial impact on model performance. Additionally, we find that training on a combination of register-specific datasets that displayed strengths in different benchmarks can lead to major improvements: registers like Opinion boosts overall model performance, while incorporating examples from register Interactive Discussion or Spoken at the expense of How-to-Instructions, Informational Description and Opinion leads to performance drops. With these findings, our study aims to open up discussion about linguistic selection of training data, and to show that register classification holds potential as a tool for analyzing pretraining data.
@@ 2 Related work
Ablation studies, where small generative models are trained with different parts of a large dataset and evaluated on downstream tasks, are a well-established method for evaluating dataset quality and optimizing data curation methods. Gao et al. (2020) validated the quality of their Pile datasets by evaluating models trained on the Pile, CommonCrawl, and CC-100. Likewise, Burchell et al. (2025) release a multilingual dataset and perform ablation studies to measure the quality of different versions of their datasets. Henriksson et al. (2025) use an LLM to line-by-line classify training documents to optimize dataset composition and evaluate with a downstream task. Longpre et al. (2024) pretrain small models on data filtered by different qualities (data age, data domain, data toxicity) and analyze the effects of this curation via model evaluation. Similarly, Sachdeva et al. (2024) use ablation-framework to find that diversity and coverage of different topics in the pretraining data are crucial to models that succeed in multiple benchmarks.
The methodology of our approach most closely resembles that of the FineWeb ablation studies (Penedo et al., 2024a). The FineWeb dataset was created by conducting analysis of tens of heuristic filters to select the data curation methods that result in the best pretraining data with respect to the downstream performance of the trained model. Currently, datasets produced by the FineWeb pipeline are available for over a thousand languages (Penedo et al., 2024b).
Current leading work on dataset composition optimization includes Su et al. (2024) and Li et al. (2024). Lee et al. (2022) established deduplication as a standard procedure in pretraining data selection pipelines. Other common approaches for selecting pretaining data from web sources include url-based filtering (e.g. Penedo et al. (2023)), text statistics such as word count, repetition, or presence of blacklisted words (e.g. Raffel et al. (2020)), filtering based on perplexity (e.g. Muennighoff et al. (2023),Weber et al. (2024)), and selection based on similarity against a selected good-quality corpus (e.g. Brown et al. (2020), Tirumala et al. (2023)). According to a survey by Albalak et al. (2024), it remains an open question whether quality filtering data improves model performance, and in which cases: Dodge et al. (2021) show that url-based pretraining data filtering leads to bias in the resulting data, while reference corpora used for similarity selection can be biased in language, topics, or demographics (Rae et al., 2022; Gururangan et al., 2022).
While register studies is a widely applied standard to analyze language variation in corpus linguistics, the term genre is more often used in the NLP field. (Sharoff et al., 2010; Petrenz & Webber, 2011; Sharoff, 2020; Kuzman et al., 2023b). However, both approaches target text categories such as opinionated, informational, and lyrical.
@@
HI-IN-HI-dtp 1/3rd of each HI-IN, HI, dtp 210 B
HI-IN-HI-dtp-OP 1/4th of each HI-IN, HI, dtp, OP 280 B
HI-IN-HI-dtp-OP-NA 1/5th of each HI-IN, HI, dtp, OP, NA 350 B
HI-IN-HI-dtp-OP-NA-ID 1/6th of each HI-IN, HI, dtp, OP, NA, ID 420 B
HI-IN-HI-dtp-OP-NA-ID-SP 1/7th of each HI-IN, HI, dtp, OP, NA, ID, SP 450 B
@@
Table 1: Datasets used in our experiments. Each dataset corresponds to one trained model. The Individual registers section describes the main register labels and hybrid & subregisters. The Combinations section contains combined register classes, and Baselines lists the baseline datasets we use in our experiments. The Available column contains the approximate token counts available on the HPLT v2 deduplicated corpus.
@@ 3.1 Data
The HPLT v2 dataset includes register information for each document, generated using the multi-label register classifier by Henriksson et al. (2024). This classifier was fine-tuned from XLM-RoBERTa-Large (Conneau et al., 2020) using the multilingual CORE corpus introduced in the same study. Although the register labels are available for multiple languages, we limit our analysis here to English due to the availability of varied and well-established benchmarks. For our experiments, we assign register labels to documents using a classification threshold of 0.4, optimized for English. The CORE register scheme is hierarchical, with 9 main registers divided into 25 subregisters that further describe the content of each document. Documents can have none, one, or multiple main registers (referred to as hybrids), along with any associated subregister labels. From these, we select all main registers, 2 subregisters, and one hybrid for our experiments. We included the subregisters and hybrids for specific reasons: the subregisters News and Description are among the largest subregister classes, each with over 100 billion tokens available—sufficient data to train models and draw meaningful comparisons with their parent registers Narrative and Informational Description. The hybrid class Instructive-Informational was included to study the differences between individual registers and hybrids, based on preliminary results showing that register-specific models trained on these classes perform well overall but differ across individual benchmark tasks. The selected register classes and their abbreviations are described with examples in Table 1. The full scheme can be found in Henriksson et al. (2024).
@@ 3.2 Model training
We replicate the training setup of Penedo et al. (2024a), using the same model architecture (Llama), model size (1.71 billion parameters), and other training parameters. In total, the models are trained up to 100 billion tokens. See Appendix A for further information on the training setup. As a baseline, we use a model trained on a random sample of the HPLT v2 deduplicated data. This allows us to compare the effects of using a single register as opposed to using data with the natural distribution of registers in the source dataset. Finally, as our setup mirrors that of Penedo et al. (2024a), we also include results for a model trained on a sample of the FineWeb dataset as a further point of comparison.
@@ 3.3 Evaluation
These benchmarks were chosen by Penedo et al. (2024a) due to their compatibility with small model size, showing reliable, low-variance and monotonic signals even with limited training data. For MMLU, we consider all 57 categories as one task and average our results accordingly. This ensures that models reaching high accuracy on MMLU but mediocre results on other benchmarks do not dominate the average results. See Appendix C for further discussion.
The selected benchmarks reflect the most common present-day expectations placed on LLMs: fluency, scientific and real-world knowledge, problem solving, and reasoning. However, this selection inevitably disadvantages certain register models. For example, models trained on the Lyrical register are likely to show poor performance, as none of the benchmarks measure poetic capabilities that would be expected from a model trained on lyrical data. This does not mean that lyrical capabilities are unwanted or unnecessary for LLMs, but rather highlights the importance of benchmark diversity and the significance of selecting benchmarks that measure abilities relevant to specific use cases.
@@ 4 Experiments
Second, we investigate the effect of further mixing registers by selecting top-performing registers from our first experiment. This allows us to observe how adding or removing specific registers affects overall model performance and whether the strengths of individual registers are preserved in combined models. The selected top-scoring register combinations are presented in Table 1 under “Combinations”. When sampling data for register combinations, we ensure that no duplicate documents are included, even when e.g. combining the hybrid class Instructive-Informational (HI-IN) with the register How-to-Instructions (HI). We chose to use the subregister Description (dtp) instead of the main-level register Informational Description (IN) because Informational Description showed signs of performance degradation in later training steps, and as noted in Section 3.1, we had to exclude some documents from the Informational Description dataset. To maintain comparability, we train all models up to 100 billion tokens, although the larger amount of combined data would allow for longer training. Our aim is not to identify optimal register combinations, but rather to explore how register-based data curation might influence model performance on different tasks.
Some of our datasets contain fewer than 100 billion tokens, which necessitates repeating data during training. This raises questions about the validity of our results, as Hoffmann et al. (2022) argue that repeating data can lead to performance degradation. However, Taylor et al. (2022) report validation loss continues to improve up to 4 epochs for a 120 billion parameter model, and Muennighoff et al. (2023) find that for the GPT-2 (Radford et al., 2019) architecture with 2-8 billion parameters, multi-epoch training has negligible effects when limited to 4 or fewer epochs. Xue et al. (2023) further demonstrate that smaller models are less susceptible to the negative effects of repeated training examples in T5-architecture models (Raffel et al., 2020). Therefore, we are reasonably confident in the results from models trained on Instructive-Informational (1.4 epochs) and Spoken (3.1 epochs). The model trained on our smallest dataset, Lyrical, performs poorly even within the first epochs, suggesting its low performance is due to register characteristics and benchmark selection rather than data repetition. None of the combination models exceed 1 epoch of training, since we sample 100 billion tokens from combinations of multiple datasets.

### v2

@@ Abstract
Pretraining data curation is a cornerstone in Large Language Model (LLM) development, leading to growing research on quality filtering of large web corpora. From statistical quality flags to LLM-based labelling systems, datasets are divided into categories, frequently reducing to a binary: those passing the filters are deemed as valuable examples, others are discarded as useless or detrimental. However, a more detailed understanding of the contribution of different kinds of texts to model performance is still largely lacking. In this article, we present the first study utilising registers or genres—a widely used standard in corpus linguistics to model linguistic variation—to curate pretraining datasets and investigate the effect of register on the performance of LLMs. We train small generative models with register classified data and evaluate them using standard benchmarks, and show that the register of pretraining data substantially affects model performance. We uncover surprising relationships between the pretraining material and the resulting models: using the News register results in subpar performance, and on the contrary, including the Opinion class, covering texts such as reviews and opinion blogs, is highly beneficial. While a model trained on the entire unfiltered dataset outperforms those trained on datasets limited to a single register, combining well-performing registers such as How-to-Instructions, Informational Description, and Opinion leads to major improvements. Furthermore, analysis of individual benchmark results reveals key differences in the strengths and drawbacks of specific register classes as pretraining data: How-to-Instructions excels at physical reasoning and sentence completion while barely crossing random baselines on world-knowledge benchmarks, while Narrative boosts performance on social interaction tasks but struggles with scientific questions. These findings show that register is an important explainer of model variation and can facilitate more deliberate and detailed future data selection practices.
@@ 1 Introduction
Recent work on scaling laws and the limits of human-generated training data available for Large Language Model (LLM) training have caused an uptick in research on pretraining data quality as opposed to quantity (Kaplan et al., 2020; Hoffmann et al., 2022; Muennighoff et al., 2023; Villalobos et al., 2024). While state-of-the-art models are still trained predominantly on large web corpora, there is increasing focus on filtering such data to remove potentially detrimental material. The selection of pretraining data often relies on simple heuristics or quality signals (Longpre et al., 2024; Ostendorff et al., 2024), sometimes together with LLM-created quality flags (Weber et al., 2024; Henriksson et al., 2025). However, language use varies along a number of dimensions other than the presumed dichotomy of positive or negative value for LLM training, ranging from persuasive, narrative, to informative texts. The effects of this variation on LLM capabilities remain unknown. In this paper, we approach pretraining data curation from a novel perspective: how do different registers (or genres) of training data affect LLM performance on commonly used benchmarks.
@@
Register studies is a widely applied paradigm in corpus linguistics that examines language variation across different situations, from casual spoken conversations to persuasive and informational settings (Biber, 1988; 1995; Biber & Conrad, 2019). Registers are defined as situationally characterised text varieties, with typical classes encompassing news, reviews and song lyrics. Register has been shown to be one of the most important predictors of linguistic variation (Biber, 2012)—and, as put by the founder of register studies, Douglas Biber, “register always matters(Biber, 2013).
@@
Web register classification has long aimed to characterise different kinds of web texts in order to better utilise and understand the web as a corpus (Kilgarriff & Grefenstette, 2003; Baroni & Kilgarriff, 2006). However, early studies focused on pre-selected sets of registers presumed to be typical of the web, and thus lacked generalizability (Santini, 2007; Sharoff et al., 2010) and subsequently failed to capture the full range of linguistic variation found on the unrestricted web. Egbert et al. (2015) addressed this gap by creating a corpus representing the full spectrum of English web registers, with Laippala et al. (2023) presenting the first deep learning-based web register classifier trained on the data and targeting the entire unrestricted web. Other studies expanded this approach to multilingual settings (Repo et al., 2021; Rönnqvist et al., 2021; Kuzman et al., 2023a; Henriksson et al., 2024), showing that web registers are identifiable in web-scale data. These advances now enable the integration of corpus linguistics frameworks with LLM pretraining data curation—our study being to the best of our knowledge the first to bridge these areas, despite the well-documented role of registers in linguistic variation.
@@
To understand the characteristics of web registers in the context of LLM pretraining, we train small generative models up to 100 billion tokens with web data with datasets curated to only contain texts from one register class. We evaluate these models using well-known benchmarks and analyse the results through average accuracy and performance on individual benchmarks, revealing how each register impacts model capabilities. Our results show that register has a substantial impact on model performance. Additionally, we find that training on a combination of register-specific datasets that displayed strengths in different benchmarks can lead to major improvements: registers like Opinion boosts model performance, while incorporating examples from register Interactive Discussion or Spoken at the expense of How-to-Instructions, Informational Description and Opinion leads to performance drops. With these findings, our study aims to open up discussion about linguistic selection of training data, and to show that register classification holds potential as a tool for analysing pretraining data.
@@ 2 Related work
Ablation studies, where small generative models are trained with different parts of a large dataset and evaluated on downstream tasks, are a well-established method for evaluating dataset quality and optimising data curation methods.
Gao et al. (2020) validated the quality of their Pile datasets by evaluating models trained on the Pile, CommonCrawl, and CC-100. Longpre et al. (2024) pretrain small models on data filtered by different qualities (data age, data domain, data toxicity) and analyse the effects of this curation via model evaluation. Similarly, Sachdeva et al. (2024) use an ablation framework to find that diversity and coverage of different topics in the pretraining data are crucial to models that succeed in multiple benchmarks. Likewise, Burchell et al. (2025) released a multilingual dataset and performed ablation studies to measure the quality of different versions of their datasets. Henriksson et al. (2025) use an LLM to line-by-line classify training documents to optimise dataset composition and evaluate with a downstream task.
The methodology of our approach most closely resembles that of the FineWeb ablation studies (Penedo et al., 2024a). The FineWeb dataset was created by conducting analyses of tens of heuristic filters to select the data curation methods that result in the best pretraining data with respect to the downstream performance of the trained model. Currently, datasets produced by the FineWeb pipeline are available for over a thousand languages (Penedo et al., 2024b).
Current leading work on dataset composition optimisation includes Su et al. (2024) and Li et al. (2024). Lee et al. (2022) established deduplication as a standard procedure in pretraining data selection pipelines. Other common approaches for selecting pretraining data from web sources include URL-based filtering (e.g. Penedo et al. (2023)), text statistics such as word count, repetition, or presence of blacklisted words (e.g. Raffel et al. (2020)), filtering based on perplexity (e.g. Muennighoff et al. (2023),Weber et al. (2024)), and selection based on similarity against a selected good-quality corpus (e.g. Brown et al. (2020), Tirumala et al. (2023)). According to a survey by Albalak et al. (2024), it remains an open question whether quality filtering data improves model performance, and in which cases: Dodge et al. (2021) show that url-based pretraining data filtering leads to bias in the resulting data, while reference corpora used for similarity selection can be biased in language, topics, or demographics (Rae et al., 2022; Gururangan et al., 2022).
While register studies is a widely applied standard to analyse language variation in corpus linguistics, the term genre is more often used in the NLP field (Sharoff et al., 2010; Petrenz & Webber, 2011; Sharoff, 2020; Kuzman et al., 2023b). However, both approaches target text categories such as opinionated, informational, and lyrical.
@@
HI-IN-HI-dtp 1/3rd sampled from HI-IN, HI, dtp
HI-IN-HI-dtp-OP 1/4th sampled from HI-IN, HI, dtp, OP
HI-IN-HI-dtp-OP-NA 1/5th sampled from HI-IN, HI, dtp, OP, NA
HI-IN-HI-dtp-OP-NA-ID 1/6th sampled from HI-IN, HI, dtp, OP, NA, ID
HI-IN-HI-dtp-OP-NA-ID-SP 1/7th sampled from HI-IN, HI, dtp, OP, NA, ID, SP
@@
Table 1: Datasets used in our experiments. Each dataset corresponds to one trained model. The Individual registers section describes the main register labels and hybrid & subregisters. The Combinations section contains combined register classes, and Baselines lists the baseline datasets we use in our experiments. The Available column contains the approximate token counts available in the HPLT v2 deduplicated corpus.
@@ 3.1 Data
The HPLT v2 dataset includes predictions of register classes as probabilities for each document, generated using the multi-label register classifier by Henriksson et al. (2024). This classifier was fine-tuned from XLM-RoBERTa-Large (Conneau et al., 2020) using the multilingual CORE corpus introduced in the same study. The CORE register scheme is hierarchical, with 9 main registers divided into 25 subregisters that further describe the content of each document. Documents can have none, one, or multiple main registers (referred to as hybrids), along with any associated subregister labels. The full register scheme can be found in Henriksson et al. (2024), and example documents are given in Appendix A.
Although the register labels are available for multiple languages, we limit our analysis to English due to the availability of varied and well-established benchmarks. For our experiments, we assign register labels to documents from the given probabilities using a classification threshold of 0.4, optimised for English. From the register classes, we select all main registers, 2 subregisters, and one hybrid for our experiments. We included the subregisters and hybrids for specific reasons: the subregisters News and Description are among the largest subregister classes, each with over 100 billion tokens available—sufficient data to train models and draw meaningful comparisons with their parent registers Narrative and Informational Description. Moreover, there has been particular interest in using news articles (e.g. de la Rosa et al. (2025)) and web encyclopedias (i.a. Wikipedia) to train LLMs, which correspond to classes News and Description in the register scheme. The hybrid class Instructive-Informational was included to study the differences between individual registers and hybrids, based on preliminary results showing that register-specific models trained on these classes perform well overall but differ across individual benchmark tasks. The selected register classes and their abbreviations are described with examples in Table 1. Subregister classes not selected for our experiments are present in the data; however, they merged to their main-level class.
@@
We also experiment with training on combinations of certain registers, which are listed in Table 1 under “Combinations”. These combination datasets are created by sampling equally from individual register datasets: for example, in the combination dataset HI-IN-HI-dtp, 1/3rd of the training samples are from Instructive-Informational (HI-IN), 1/3rd from How-to-Instructions (HI), and 1/3rd from Description (dtp), with our dataset creation step ensuring no duplicates are present in the combined data. See Section 4 for further information and motivation for these datasets.
@@ 3.2 Model training
We replicate the training setup of Penedo et al. (2024a), using the same model architecture (Llama), model size (1.71 billion parameters), and other training parameters. In total, the models are trained up to 100 billion tokens. See Appendix B for further information on the training setup. As a baseline, we use a model trained on a random sample of the HPLT v2 deduplicated data. This allows us to compare the effects of using a single register as opposed to using data with the natural distribution of registers in the source dataset. Finally, as our setup mirrors that of Penedo et al. (2024a), we also include results for a model trained on a sample of the FineWeb dataset as a further point of comparison and as a reference for interpreting our findings.
@@ 3.3 Evaluation
These benchmarks were chosen by Penedo et al. (2024a) due to their compatibility with small model size, showing reliable, low-variance and monotonic signals even with limited training data, which is crucial for our experiments. For MMLU, we consider all 57 categories as one task and average our results accordingly. This ensures that models reaching high accuracy on MMLU but mediocre results on other benchmarks do not dominate the average results. See Appendix D for further discussion.
The selected benchmarks reflect the most common present-day expectations placed on LLMs: fluency, scientific and real-world knowledge, problem solving, and reasoning. However, this selection inevitably disadvantages certain register models. For example, models trained on the Lyrical register are likely to show poor performance, as none of the benchmarks measure poetic capabilities that would be expected from a model trained on lyrical data. This does not mean that lyrical capabilities are unwanted or unnecessary for LLMs, but rather highlights the importance of benchmark diversity and the significance of selecting benchmarks that measure abilities relevant to specific use cases. In our case, although the benchmark selection disadvantages some registers and their expected performance more than others, using the full range of registers covering all kinds of web texts yields a complete picture of registers’ effects on LLM pretraining.
@@ 4 Experiments
Second, we investigate the effect of further mixing registers by selecting top-performing registers from our first experiment. This allows us to observe how adding or removing specific registers affects overall model performance and whether the strengths of individual registers are preserved in combined models. The selected top-scoring register combinations are presented in Table 1 under “Combinations”. When sampling data for register combinations, we ensure that no duplicate documents are included, even when combining the hybrid class Instructive-Informational (HI-IN) with the register How-to-Instructions (HI). We chose to use the subregister Description (dtp) instead of the main-level register Informational Description (IN) because Informational Description showed signs of performance degradation in later training steps, and as noted in Section 3.1, we had to exclude some documents from the Informational Description dataset. To maintain comparability, we train all models up to 100 billion tokens, although the larger amount of combined data would allow for longer training. Our aim is not to identify optimal register combinations, but rather to explore how register-based data curation might influence model performance on different tasks.
Some of our datasets contain fewer than 100 billion tokens, which necessitates repeating data during training. This raises questions about the validity of our results, as Hoffmann et al. (2022) argue that repeating data can lead to performance degradation. However, Taylor et al. (2022) report validation loss continues to improve up to 4 epochs for a 120 billion parameter model, and Muennighoff et al. (2023) find that for the GPT-2 (Radford et al., 2019) architecture with 2-8 billion parameters, multi-epoch training has negligible effects when limited to 4 or fewer epochs. Xue et al. (2023) further demonstrate that smaller models are less susceptible to the negative effects of repeated training examples in T5-architecture models (Raffel et al., 2020). Therefore, we are reasonably confident in the results from models trained on Instructive-Informational (1.4 epochs) and Spoken (3.1 epochs). The model trained on our smallest dataset, Lyrical (5 epochs), performs poorly even within the first epochs, suggesting its low performance is due to register characteristics and benchmark selection rather than data repetition. None of the combination models exceed 1 epoch of training, since we sample the 100 billion tokens from combinations of multiple datasets, together exceeding 100 billion tokens.
We additionally compare our register scheme to three other quality classifiers, FineWeb-edu classifier (Penedo et al., 2024a), the DCLM quality classifier (Li et al., 2024), and NVIDIA’s NemoCurator Quality Classifier DeBERTa (He et al., 2023). These experiments and their results are presented in Appendix C.